A cylinder of cooking gas supplied by Indian Oil Corporation is assumed to contain 14 kg of butane (ΔHc for C4H10 = - 2600 kJ mol-1). If a small family of three persons, requires 10,000 J of heat energy per day for cooking, the gas in the cylinder would last for

Grade:12th pass

3 Answers

Ashish bedi
7 years ago
14kg butane =241.3 molesTotal energy released =241.3*2600*1000 joules. Days gas last = 241.3*2600*1000/1000 0= 62738 days
Deep Mukherjee
7 years ago
63 days= 62.7=627380/10000 Days gas last = 627380.=241.3*2600KJTotal energy released =241.3 moles14kg butane =
 
Deep Mukherjee
7 years ago
 14kg butane =241.3 moles.
Total energy released =241.3*2600KJ = 627380.
Days gas last = 627380/10000=62.7= 63 days.